### 01 Apr 2026 v2

# ARCS: Autoregressive Circuit Synthesis with Topology-Aware Graph Attention and Spec Conditioning

Abstract
This paper presents ARCS (Autoregressive Circuit Synthesis), a system for amortized analog circuit generation that produces complete, SPICE-simulatable designs (topology and component values) in milliseconds rather than the minutes required by search-based methods. A hybrid pipeline combining two learned generators (a graph VAE and a flow-matching model) with SPICE-based ranking achieves 99.9% simulation validity (reward 6.43/8.0) across 32 topologies using only 8 SPICE evaluations, 40 fewer than genetic algorithms. For single-model inference, a topology-aware Graph Transformer with Best-of-3 candidate selection reaches 85% simulation validity in 97 ms, over 600 faster than random search. The key technical contribution adapts Group Relative Policy Optimization (GRPO) to multi-topology circuit reinforcement learning, resolving a critical failure mode of REINFORCE (cross-topology reward distribution mismatch) through per-topology advantage normalization. This improves simulation validity by +9.6 percentage points over REINFORCE in only 500 RL steps ( fewer). Grammar-constrained decoding additionally guarantees 100% structural validity by construction via topology-aware token masking.
I Introduction
Designing an analog circuit takes days. Engineers select a topology, choose component values, simulate, and iterate, a cycle repeated until specifications are met. Recent machine learning approaches automate parts of this workflow, but each leaves critical gaps.
Topology-only generation. AnalogGenie [1] generates circuit topologies using an autoregressive transformer over pin-level Eulerian sequences, achieving 93.2% validity after PPO fine-tuning. However, it produces no component values; a separate genetic algorithm must size every generated circuit, adding minutes of computation. It also lacks specification conditioning: circuits are generated randomly, then filtered post-hoc.
Text-based generation. CircuitSynth [2] and AutoCircuit-RL [3] fine-tune large language models (GPT-Neo, 2.7B parameters) on raw SPICE netlist text. These models predict characters, not circuit components, leading to syntax errors in generated netlists and no semantic understanding of component values.
This work. ARCS (Figure 1) addresses all three gaps simultaneously. Given one of 32 parameterized topology templates and a target specification, ARCS generates complete component-value assignments in a single forward pass. The three core contributions are:
1.
Group Relative Policy Optimization (GRPO) for multi-topology RL: REINFORCE suffers from cross-topology reward distribution mismatch: easy topologies dominate gradient updates while hard topologies are abandoned. GRPO resolves this via per-topology advantage normalization (Eq. 6), achieving 53.13.1% simulation validity (+9.6 percentage points (pp) over REINFORCE) with 10 fewer RL steps.
2.
Grammar-constrained decoding: A state-machine-based token masking scheme guarantees 100% structural validity by construction, without RL training or post-hoc filtering. Combined with Best-of-3 candidate selection, this yields 85% simulation validity in 97 ms.
3.
Hybrid multi-source ranking: A pipeline combining two complementary generators (a graph-structured VAE (VCG) and a Constrained Circuit Flow Matching model (CCFM)) with SPICE-based ranking achieves 99.9% simulation validity and reward 6.43/8.0 with only 8 SPICE evaluations (40 fewer than GA).
The key contribution of ARCS is amortized inference: a trained model that produces reasonable first-shot designs in 20 ms, enabling rapid prototyping and design-space exploration at a cost 1000 lower than conventional search. ARCS does not yet match the per-design quality of search baselines (5.48 vs. 7.48 reward), but the paradigms are complementary: ARCS-seeded GA recovers 96.6% of cold-start quality with 49% fewer simulations.
II Related Work
II-A Autoregressive Circuit Generation
AnalogGenie [1] introduced Eulerian circuit representations for autoregressive generation, training a GPT decoder (11.8M parameters) on 3,502 circuits manually curated from IEEE publications. After PPO fine-tuning, it achieves 93.2% structural validity. Key limitations: no component values in the vocabulary, no specification conditioning, and reliance on a GA for post-hoc sizing.
CircuitSynth [2] and AutoCircuit-RL [3] fine-tune GPT-Neo (2.7B parameters) on SPICE netlist text, treating circuit generation as a text completion task. While this includes component values in the output, the model operates at the character level with no circuit-specific inductive bias.
LaMAGIC [7] generates analog IC topologies using language models with graph-based representations. While it introduces circuit-aware tokenization, it targets transistor-level IC design rather than board-level circuits with discrete component values. AnalogXpert [15] incorporates circuit design expertise into large language models for topology synthesis, demonstrating that domain knowledge can guide LLM-based generation. AstRL [14] applies deep reinforcement learning to analog and mixed-signal circuit synthesis, training an RL agent to select topologies and size components end-to-end. INSIGHT [16] uses an autoregressive transformer as a fast surrogate for SPICE simulation, achieving near-simulator accuracy at orders-of-magnitude lower cost; however, it is a predictor of circuit performance, not a generator of new designs.
II-B Circuit Sizing and Optimization
AutoCkt [6] uses deep reinforcement learning for analog circuit sizing, treating component values as a continuous action space. Wang et al. [23] combine graph neural networks with RL for transferable transistor sizing across circuit topologies. Lyu et al. [22] apply Bayesian optimization to analog sizing, achieving efficient exploration with Gaussian process surrogate models. CktGNN [5] uses graph neural networks for topology generation but is limited to operational amplifiers. Fan et al. [17] propose a graph-transformer surrogate model for power converter topologies, achieving strong prediction accuracy but focusing on evaluation rather than generation. FALCON [4] combines GNN-based prediction with layout co-optimization but is not generative for new topologies. DiffCkt [13] applies diffusion models to transistor-level analog circuit generation via a hybrid neural network framework, demonstrating that generative models beyond VAEs and autoregressive approaches can produce valid circuit topologies.
II-C Flow-Based Generative Models
Conditional flow matching (CFM) [18] learns continuous-time transport maps between a prior and data distribution via optimal-transport velocity fields, offering straighter sampling trajectories than diffusion models and avoiding the KL-constrained latent space of VAEs. CFM has been applied to molecular generation [24] and protein design but not, to date, to circuit synthesis. For discrete graph generation, DiGress [21] applies discrete denoising diffusion to molecule and planar graphs; however, it does not handle mixed discrete-continuous outputs (topology + values) or specification conditioning. ARCS introduces CCFM (Section III-F), the first application of flow matching to analog circuit generation.
II-D Key Differences
| Values | Specs | Graph | Valid | Speed | Data | |
| AnalogGenie | × | × | × | 93.2% | 1 s | Manual |
| LaMAGIC | × | × | ✓ | N/A | N/A | Synthetic |
| CircuitSynth | ✓ | × | × | N/A | 10 s | RL only |
| AstRL | ✓ | ✓ | × | N/A | N/A | RL only |
| DiffCkt | ✓ | × | ✓ | N/A | N/A | Synthetic |
| ARCS | ✓ | ✓ | ✓ | 100%† | 0.02 s | Auto |
| †Structural validity with grammar-constrained decoding. | ||||||
| Sim-validity: 85% (Best-of-3), 99.9% (hybrid pipeline). | ||||||
ARCS is the first system that jointly generates component values conditioned on both topology and target specifications while incorporating circuit-graph inductive bias (Table I). Notably, ARCS performs conditioned component sizing across 32 predefined topology templates, not topology discovery. The topology is selected from a known library; the model’s task is to predict all component values that satisfy the target specification. AnalogGenie generates diverse topologies but outputs only graph structure without component values. ARCS produces immediately simulatable netlists, closing the gap between generation and verification.
III Method
III-A Problem Formulation
Given a target specification (e.g., V, A) and a topology label (e.g., “buck”), generate a sequence of component tokens where each pair specifies a component type and its value. The generated circuit should be electrically valid and meet the target specification when simulated.
III-B Tokenizer
The tokenizer defines a domain-specific vocabulary of 706 tokens organized into seven semantic categories:
Special tokens (5): START, END, PAD, SEP, INVALID.
Component tokens (20): MOSFET_N, RESISTOR, CAPACITOR, INDUCTOR, DIODE, OPAMP, TRANSFORMER, etc.
Topology tokens (40): Slots for 34 named circuit topologies plus 6 reserved tokens for future expansion.
Spec tokens (20): SPEC_VIN, SPEC_VOUT, SPEC_IOUT, SPEC_FSW, SPEC_GAIN, SPEC_BW, etc.
Pin tokens (21): Component pin labels (PIN_DRAIN, PIN_POS, etc.).
Net/connection tokens (100): Circuit net identifiers.
Value tokens (500): Log-uniformly spaced bins covering to , providing 28 bins per decade for sub-decade resolution.
Each circuit is encoded as a flat sequence:
The log-uniform binning ensures that physically meaningful value ratios (e.g., vs. ) map to proportional token distances, and data augmentation via random component order shuffling () teaches the model that component ordering is arbitrary.
| Example: Generated Buck Converter |
| START TOPO_BUCK SEP |
| SPEC_VIN VAL_12.0 SPEC_VOUT VAL_5.0 |
| SPEC_IOUT VAL_1.0 SEP |
| INDUCTOR VAL_100H CAPACITOR VAL_22F |
| RESISTOR VAL_5m MOSFET VAL_15m |
| END |
| Decoded to SPICE netlist |
| L1 sw_node vout 1.000e-04 |
| C1 cap_node 0 2.200e-05 IC=5.0 |
| S1 input sw_node pwm 0 SMOD |
| Rload vout 0 5.0 |
III-C Model Architecture
ARCS explores three progressively more expressive architectures, all sharing a common GPT-style backbone: , , , , SwiGLU activations [8], RMSNorm [9], and learned token-type embeddings (7 categories). All models use weight-tied language model heads.
(A) Baseline GPT (6.5M parameters). A standard decoder-only transformer with a single weight-tied output head over the full vocabulary. This model treats all tokens (structural and value) identically.
(B) Two-Head Model (6.8M parameters; +4.7%). Predicting which component to place and predicting what value it should have are fundamentally different tasks. The two-head model separates these:
A structure head (weight-tied with the token embedding) predicts component type, topology, spec, and special tokens.
A value head (dedicated 2-layer SiLU MLP, with residual connection, followed by a linear projection to ) specializes in predicting numerical value tokens.
At each timestep, the model selects which head to use based on whether the next token is expected to be a value token (determined by the preceding component token).
(C) Graph Transformer (6.8M parameters; +5.0%). Circuit components form a graph defined by their electrical connectivity. The Graph Transformer injects this structure via two mechanisms:
1.
Topology-aware attention bias. For each of the 32 topologies, ARCS precomputes a binary adjacency matrix where is the maximum number of components. During self-attention, a learned scalar bias augments the attention logits for positions where :
(1) This gives the model a soft prior that electrically connected components should attend to each other more strongly.
2.
Random-walk positional encoding (RWPE). Each component receives a -dimensional positional feature derived from the topology’s random-walk transition matrix , where is the degree matrix. For node , the RWPE is the vector of return probabilities after steps:
(2) with walk lengths. These features are precomputed per topology and projected to via a learned two-layer MLP (, GELU activation). The following value token shares the same RWPE as its parent component. Unlike ordinal position indices, RWPE encodes each component’s structural role in the circuit graph; for example, high-degree hub nodes (like the switch node in a buck converter) have higher return probabilities than peripheral nodes (like the load resistor).
The Graph Transformer retains the two-head output structure from model (B), adding 17,648 parameters total: 17,216 for the RWPE projection MLP and 432 for graph attention biases (per-head adjacency scalars and edge-type embeddings).
III-D Automated Data Generation
Unlike AnalogGenie’s manually curated dataset (3,502 circuits from IEEE papers), ARCS uses a fully automated pipeline:
1.
Template definition: Parameterized SPICE netlist templates with physical component bounds (E-series snapped) for each of 32 topologies.
2.
Random sampling: Component values drawn log-uniformly within bounds.
3.
SPICE simulation: ngspice transient/AC simulation with automatic metric extraction (efficiency, output accuracy, gain, bandwidth, etc.).
4.
Filtering: Both valid and invalid designs are stored; the model learns to avoid failure modes.
This pipeline generated 62,000 circuits (41,064 valid after simulation) across 34 topologies. After augmentation via component-order shuffling, the training set contains 205,000 sequences.
III-E Training
Stage 1: Supervised pre-training. Next-token prediction with cross-entropy loss, applying weight to value tokens to compensate for their larger vocabulary:
where for value tokens and otherwise. Training: 100 epochs, batch size 64, AdamW (, ), cosine LR schedule ( peak) with 5-epoch linear warmup.
Stage 2: REINFORCE [25] with SPICE reward. Stage 2 fine-tunes the best supervised checkpoint with policy gradient optimization, using SPICE simulation as an oracle reward signal:
where is an exponential moving average baseline and is an adaptive KL coefficient [3] with . The reward function (max 8.0) decomposes as:
The sub-rewards are domain-specific and spec-aware: power converters use output voltage error () and efficiency; amplifiers use gain accuracy vs. target () and bandwidth; filters use cutoff frequency accuracy () and passband gain; oscillators use frequency accuracy () and amplitude. This ensures the reward measures how well generated circuits meet their target specifications, not just functional correctness. RL training runs for 5,000 steps (batch size 8, lr , temperature 0.8, top- 50), requiring 15 hours on an Apple M3 with MPS acceleration.
Stage 3: Group Relative Policy Optimization (GRPO). Stage 2 reveals a critical failure mode: global-baseline REINFORCE suffers from cross-topology reward distribution mismatch. Power converters have max reward 8.0 while amplifiers reach 4.0; when the global baseline settles near the cross-topology mean, easy topologies receive systematically positive advantages while hard topologies receive negative ones, causing the policy to abandon difficult circuits entirely (Buck: 100%10% sim-valid under REINFORCE).
GRPO resolves this by computing advantages within topology groups, inspired by the group-relative advantage estimation of Shao et al. [20] but applied per-topology rather than per-prompt. Each RL step samples topologies and generates circuits per topology (a “group”). Advantages are z-scored per group:
where are the mean and standard deviation of rewards within topology ’s group. This ensures that every topology contributes meaningful gradient signal regardless of its absolute reward scale. The policy gradient becomes:
The training uses topologies and circuits per step. With only 500 GRPO steps (1 hour, 6,000 SPICE simulations), the model achieves 53.13.1% simulation validity, surpassing both supervised learning (45.43.0%) and 5,000-step REINFORCE (43.51.2%), validating the hypothesis that per-topology normalization is essential for multi-topology RL.
III-F Constrained Circuit Flow Matching (CCFM)
While the autoregressive approach generates circuits token-by-token, ARCS additionally includes a non-autoregressive generative model that generates all circuit components in parallel via conditional flow matching [18].
Architecture. CCFM operates in the latent space of a pre-trained ValidCircuitGen (VCG) encoder, a graph-structured VAE (4.0M parameters) trained on 41,064 valid circuits with 5 differentiable structural constraints (type coherence, adjacency symmetry, graph connectivity, degree bounds, value range). The flow network (3.7M parameters) consists of 4 DiT-style transformer blocks [19] with adaptive layer normalization conditioned on both time embedding and specification embedding :
where is a sinusoidal time embedding and is a learned spec projection. Each DiT block applies adaptive LayerNorm followed by multi-head self-attention and a SwiGLU feed-forward network, with all conditioning applied via scale-and-shift modulation.
Training. CCFM uses optimal-transport conditional flow matching [18]. Given paired samples and from the VCG latent space, the interpolant is and the target velocity is . The loss is:
The VCG encoder is frozen during flow matching training. After 100 epochs (25 s/epoch on MPS), CCFM achieves validation loss 0.14 (from 0.68 initial).
Constraint-guided sampling. During ODE integration (Euler method, 50 steps), CCFM projects the velocity toward the feasible circuit set using learned per-constraint guidance weights , initialized to zero and trained end-to-end:
where applies stronger guidance early (when samples are noisy) and relaxes as the trajectory approaches the data distribution. The five constraints match VCG’s: type coherence, adjacency symmetry, connectivity, degree bounds, and value range.
III-G Grammar-Constrained Decoding
A fundamental limitation of unconstrained autoregressive sampling is that the model may generate structurally invalid token sequences (e.g., two consecutive component tokens without an intervening value, or a wrong component type for the target topology). While RL can improve structural validity to 100% (Section V), it does so statistically, without formal guarantees.
ARCS introduces grammar-constrained decoding, which applies a topology-aware token mask at each autoregressive step, restricting the sampling distribution to only structurally valid continuations. This provides 100% structural validity by construction, without RL training or post-hoc filtering.
Grammar state machine. The decoder maintains a finite-state machine with two alternating phases: EXPECT_COMP (the next token must be a valid component type or the END token) and EXPECT_VAL (the next token must be a value token). The state machine transitions deterministically based on the last emitted token. Figure 3 illustrates the state transitions.
Constraint levels. ARCS defines three levels of increasing strictness:
1.
Grammar: Enforces the component–value alternation pattern. At EXPECT_COMP steps, only component and special tokens are allowed; at EXPECT_VAL steps, only value tokens.
2.
Topology: Extends Grammar by restricting the set of allowed component types to exactly those required by the target topology (e.g., a buck converter requires INDUCTOR, CAPACITOR, DIODE, MOSFET_N). This ensures correct component types and counts.
3.
Full: Extends Topology by restricting value tokens to the physically valid range for each component type (e.g., inductors in [1 H, 10 mH], capacitors in [1 pF, 10 mF]).
Masked sampling. At each step , the decoder applies a binary mask over the vocabulary before the softmax:
where are the raw logits and zeroes out invalid tokens. Since the mask is a deterministic function of the grammar state and topology, it adds negligible computational overhead (pre-computed token sets, 1 ms per sequence). In practice, constrained decoding is actually faster than unconstrained sampling because the model converges to valid END tokens more quickly, eliminating wasted tokens from invalid continuations.
Lagrangian constraint loss. For training with constraint awareness, the model additionally optimizes a differentiable Lagrangian loss:
where are constraint violation penalties (structural validity, component correctness, value range adherence) and are adaptive Lagrange multipliers updated via dual gradient ascent. This allows training to internalize the constraints, improving generation quality even when decoding without masks.
III-H Learned Reward Model
To improve Best-of- candidate ranking beyond model confidence, ARCS also trains a learned reward model, a bidirectional transformer encoder that predicts SPICE simulation reward from token sequences, following the verifier approach of Cobbe et al. [12]. The architecture uses 2 encoder layers with 128-dim hidden states, 4 attention heads, and SwiGLU FFN, followed by mean-pooling and a 2-layer MLP producing a scalar reward prediction. Token embeddings are warm-started from the generator via SVD-based projection when the embedding dimensions differ. The 666K-parameter model is trained on 41K circuits from the automated data pipeline using Huber loss.
IV Experimental Setup
IV-A Circuit Topologies
ARCS defines 34 circuit topologies across three tiers, of which 32 are evaluated in the main experiments:
Tier 1 (7 power converters; 5 evaluated): Buck, Boost, Buck-Boost, Cuk, SEPIC.
Tier 2 (9 signal circuits): Inverting/non-inverting/ instrumentation/differential amplifiers, Sallen-Key LP/HP/BP filters, Wien bridge and Colpitts oscillators.
Tier 2b (18 extended circuits): BJT amplifiers (common emitter/collector/base, cascode), current sources, regulators, additional oscillators (Hartley, phase shift), filters (state variable, twin-T notch), and extended power (half bridge, push pull, charge pump, voltage doubler, zeta converter).
IV-B Baselines
Random Search (RS): For each target spec, uniformly sample 200 parameter vectors (log-uniform within bounds), simulate each, and return the highest-reward design. Total: 200 SPICE simulations per spec.
Genetic Algorithm (GA): BLX- crossover in log-space, Gaussian mutation, tournament selection. Table II uses population 30 over 20 generations (630 evals). The statistical evaluation (Table IV) uses 20 repeats with population 16 (320 evals per spec) to match the compute budget of the hybrid pipeline.
Both baselines have a fundamental advantage: they search directly in the known parameter space by evaluating hundreds of value combinations, while ARCS produces values in a single forward pass from a spec prefix. Note that ARCS is given the correct topology token in its input prefix (making it a topology-conditioned generation task, not topology discovery), so the information asymmetry lies in the search budget, not topology access. This makes the comparison meaningful: given the same topology, can a learned model’s single-shot prediction compete with iterative search?
IV-C Evaluation Protocol
Each method is evaluated on 160 conditioned test specs (10 per topology). Metrics:
Structural validity: Well-formed token sequence that decodes to a valid circuit.
Sim. success: SPICE convergence rate.
Sim. validity: Physically plausible simulation results (no negative efficiency, reasonable output).
Reward: Domain-specific score (max 8.0), combining accuracy, efficiency, and quality metrics.
Wall time: End-to-end time per design including SPICE simulation for baselines.
V Results
V-A Architecture Comparison and Main Results
| Method | Params | Struct | SimValid | Reward |
| Random Search | — | — | 81.2% | 7.28 |
| Genetic Alg. | — | — | 80.0% | 7.48 |
| Baseline GPT (SL) | 6.5M | 86.01.9% | 40.10.7% | 3.370.06 |
| Two-Head (SL) | 6.8M | 96.21.0% | 50.42.2% | 3.890.09 |
| Graph Trans. (SL) | 6.8M | 93.21.4% | 45.43.0% | 3.860.15 |
| REINFORCE (5000) | 6.5M | 95.51.3% | 43.51.2% | 3.740.04 |
| GT+GRPO (500) | 6.8M | 96.60.5% | 53.13.1% | 4.150.08 |
| GT+GRPO (3500) | 6.8M | 92.91.3% | 48.82.2% | 3.790.05 |
| VCG (graph VAE) | 4.0M | 100% valid | — | |
| CCFM (flow match) | 7.7M | 100% valid | — | |
Table II shows the unified comparison averaged over 5 random seeds. Architecture matters. The two-head architecture improves simulation validity by pp over the baseline by decoupling structural and numerical prediction. The Graph Transformer achieves pp over baseline via topology-aware attention biases (Eq. 1). Notably, the two-head model (50.4%) outperforms the Graph Transformer under supervised learning alone (45.4%): the GT’s relational inductive bias is not fully exploited by cross-entropy training on fixed data, but becomes decisive when combined with GRPO reinforcement learning, where GT+GRPO (53.1%) surpasses both.
GRPO resolves the RL regression. REINFORCE slightly degrades simulation validity ( pp vs. SL) due to cross-topology reward mismatch, while GRPO improves it by pp over GT SL and pp over REINFORCE. These differences exceed the 3 pp standard deviations, confirming significance across seeds. GRPO uses only 500 RL steps (6,000 SPICE simulations, 1 hour) compared to REINFORCE’s 5,000 steps (40,000 simulations, 15 hours) , a 10 reduction in training cost. Extended GRPO training (3,500 total steps) shows diminishing returns: sim-validity drops to 48.82.2%, suggesting that 500 steps is the optimal early-stopping point for this model scale. This confirms the hypothesis (Section III-E) that per-topology advantage normalization is essential for multi-topology RL.
Non-autoregressive alternatives. VCG and CCFM both achieve 100% structural validity via differentiable constraints, reaching 85.3% simulation validity (116/136) across 34 topologies. CCFM trains 3 faster than VCG (25 s vs. 71 s per epoch) while matching validity.
V-B Hybrid Multi-Source Ranking
| Method | Struct | SimSuccess | SimValid | Reward |
| VCG-only | 100% | 95.6% | 85.3% | 5.53 |
| CCFM-only | 100% | 97.1% | 85.3% | 5.50 |
| Hybrid (VCG+CCFM) | 100% | 97.1% | 94.1% | 6.37 |
Multi-source ranking is strongly complementary: selecting by SPICE reward improves average reward by +0.84 over VCG-only and +0.87 over CCFM-only.
V-C Statistical Evaluation with Baselines
| Method | SPICE Evals | Reward (95% CI) | SimValid |
| Random Search | 1 | 5.18 [5.05, 5.31] | 91.4% |
| Genetic Alg. | 320 | 7.56 [7.53, 7.60] | 100.0% |
| VCG only | 1 | 5.34 [5.27, 5.42] | 94.0% |
| CCFM only | 1 | 5.51 [5.44, 5.58] | 95.7% |
| Hybrid | 8 | 6.43 [6.38, 6.48] | 99.9% |
At equal compute (1 SPICE evaluation), both VCG and CCFM outperform random sampling. The hybrid pipeline with 8 candidates achieves 99.9% simulation validity and reward 6.43 using 40 fewer SPICE evaluations than GA (7.56). Each pipeline component contributes significant improvements (): VCG (5.34) +CCFM (5.51) +hybrid ranking (6.43).
V-D Ablation Studies
| Variant | Struct | SimValid | Reward |
| ARCS + RL (full) | 98.1% | 52.5% | 3.49 |
| No RL (supervised only) | 90.6% | 46.9% | 3.24 |
| No spec conditioning | 58.8% | 38.8% | 2.60 |
| Tier 1 only (7 topos) | 100% | 20.6% | 3.77 |
Removing spec conditioning drops reward by and structural validity to 58.8%, confirming that the specification prefix is essential. RL improves reward by and validity by pp. The Tier 1-only model achieves the highest per-design reward (3.77) but covers only 7 topologies; 32 topologies with lower average validity is the more useful configuration.
Signal circuits show the clearest GRPO advantage (663% sim-validity vs. 503% REINFORCE). Power converters remain hardest due to switching dynamics; GRPO still outperforms alternatives. Full per-topology results are in Appendix Table XI.
V-E Warm-Start Experiment
A natural question is whether ARCS’s fast inference can complement search-based optimization. The warm-start strategy works as follows: ARCS generates 3 candidate designs, and the best seeds a GA population of 20 individuals for 5 generations (113 SPICE simulations). The comparison is cold-start GA (20 individuals, 10 generations, 220 simulations).
| Method | Reward | Sims | Time |
| ARCS only | 5.45 | 3 | 0.8 s |
| GA cold-start | 7.43 | 220 | 47.3 s |
| GA warm-start | 7.18 | 113 | 19.7 s |
Table VI shows that on the 14 topologies where ARCS produces valid initial designs (Wien bridge and Colpitts fall back to cold-start), GA-Warm achieves 96.6% of cold-start quality (7.18 vs. 7.43 reward) while using only 113 simulations, a 49% reduction. For signal-processing topologies, warm-start matches cold-start exactly. On boost converter, warm-start outperforms cold-start (7.95 vs. 7.84), suggesting that ARCS’s learned initialization lands in a better basin of attraction. The 58% wall-clock speedup validates ARCS as a practical initializer for search-based design workflows.
V-F Constrained Decoding
This section evaluates the grammar-constrained decoding scheme (Section III-G) across three constraint levels (Grammar, Topology, Full) plus an unconstrained baseline (None). To isolate constraints from model quality, the evaluation uses a randomly initialized model (no training), demonstrating that the constraints alone suffice for structural correctness.
| Level | Struct % | Comp % | Avg Comp | Time |
| None | 0.0% | 0.0% | — | 269 ms |
| Grammar | 100% | 0.0% | 18.6 | 125 ms |
| Topology | 100% | 100% | 4.3 | 27 ms |
| Full | 100% | 100% | 4.3 | 25 ms |
All three constraint levels achieve 100% structural validity, compared to 0% unconstrained, a formal guarantee across all 32 topologies. Topology constraints additionally ensure 100% component type correctness. Counter-intuitively, constrained decoding is faster (25 ms vs. 269 ms unconstrained) because topology constraints terminate sequences promptly instead of generating meandering token streams. Since constraints are orthogonal to model quality (working even with random initialization), they compose freely with RL fine-tuning.
V-G Inference-Time Compute Scaling
ARCS generates a circuit in 30 ms. Generating candidates and selecting the best, without any SPICE simulation, trades minimal additional inference cost for improved quality. The ranker uses mean log-probability (model confidence), which is computed for free during autoregressive sampling. This is analogous to recent “test-time compute scaling” [11] results in language modeling, where generating more candidates and selecting by a scoring function yields substantial quality gains at low marginal cost.
| Conf. | SimValid | Reward | Time | Speedup | |
| 1 | 1.16 | 72.5% | 5.01 | 35 ms | 1,680 |
| 3 | 0.82 | 85.0% | 5.48 | 97 ms | 606 |
| 5 | 0.74 | 85.0% | 5.40 | 164 ms | 359 |
| 10 | 0.67 | 77.5% | 5.14 | 324 ms | 181 |
| 20 | 0.57 | 82.5% | 5.18 | 641 ms | 92 |
| 50 | 0.52 | 82.5% | 5.10 | 1.6 s | 37 |
| RS | — | 81.2% | 7.28 | 58.8 s | 1 |
SPICE reward peaks at (5.48, +9.3% over single-shot) then plateaus, even as model confidence continues improving, indicating that log-probability is a useful but imperfect proxy for simulation quality. Best-of-3 at 97 ms remains 600 faster than random search (58.8 s).
The non-monotonic reward curve at motivates the learned reward model (Section V-H) for more effective ranking.
V-H Learned Reward Model
To address the confidence–reward misalignment (Section V-G), a 666K-parameter reward model (2-layer bidirectional transformer encoder, 128-dim, 4 heads) is trained on 41K SPICE-simulated circuits using Huber loss. Token embeddings are warm-started from the generator via SVD projection.
| Metric | Value |
| Pearson correlation () | 0.988 |
| Spearman rank corr. () | 0.947 |
| MAE | 0.147 |
| RMSE | 0.390 |
| Within 0.5 | 93.4% |
| Within 1.0 | 97.8% |
The reward model achieves correlation with ground-truth reward (Table IX).
Ranking comparison. A comparison of confidence-based and reward-model-based ranking for Best-of- candidate selection, evaluating the selected circuit via SPICE simulation (16 test specs, all topologies).
| SL Generator | RL Generator | |||||
| Conf | RM | % | Conf | RM | % | |
| 1 | 5.16 | 5.16 | — | 5.06 | 5.06 | — |
| 3 | 5.05 | 5.15 | +2.0 | 5.08 | 5.18 | +2.1 |
| 5 | 5.03 | 4.71 | 6.5 | 5.12 | 5.16 | +0.7 |
| 10 | 5.08 | 5.22 | +2.7 | 5.30 | 5.16 | 2.5 |
| 20 | 5.18 | 4.81 | 7.2 | 5.04 | 5.17 | +2.5 |
At , the reward model consistently outperforms confidence ranking for both SL and RL generators ( and ), confirming that the learned reward signal captures quality aspects that log-probability misses. The reward model is most reliably valuable at moderate (3), providing a consistent, training-free quality boost at negligible cost.
VI Discussion
Architecture vs. RL vs. GRPO. Architectural changes (BaselineGraph Transformer) yield reward from inductive bias alone. REINFORCE degrades validity ( pp) due to cross-topology reward mismatch, while GRPO improves it ( pp over SL) with 10 fewer steps. Per-topology advantage normalization is essential when training spans circuit families with heterogeneous reward scales.
Amortized vs. iterative design. ARCS and search baselines are complementary: ARCS-seeded GA recovers 96.6% of cold-start quality with 49% fewer simulations (Section V-E).
Structure by grammar, semantics by learning. Constrained decoding guarantees 100% structural validity by construction; GRPO optimizes value quality within the valid subspace. This decomposition mirrors the separation of syntax and semantics in programming language design [10].
Scalability. ARCS scales to new topologies by writing a single SPICE template, with no manual curation needed. Extending to transistor-level IC circuits requires more templates and likely larger models.
Limitations.
Per-design quality remains below search baselines (5.48 Best-of-3 vs. 7.48 GA). Scaling to 50–100M parameters with more training data is the primary direction for improvement.
While constrained decoding guarantees structural validity, simulation validity (i.e., whether SPICE produces correct output) still depends on learned component values.
The value tokenizer’s 500-bin resolution (28 bins/decade) limits precision to 3.5% relative error, which may be insufficient for sensitive RF or precision analog circuits.
Future work. Scaling to 50–100M parameters, curriculum learning over topology difficulty, end-to-end CCFM fine-tuning with SPICE-in-the-loop, and extending constraints to enforce Kirchhoff’s laws.
Code and data availability. Code, data, and trained checkpoints are available at https://github.com/tusharpathaknyu/ARCS.
VII Conclusion
ARCS generates complete, SPICE-simulatable analog circuits conditioned on target specifications across 32 topologies, in milliseconds rather than minutes. Three findings stand out. First, per-topology advantage normalization (GRPO) resolves the cross-topology reward mismatch that causes REINFORCE to regress, improving simulation validity by +9.6 pp with 10 fewer RL steps. Second, grammar-constrained decoding guarantees 100% structural validity by construction, reaching 85% simulation validity in 97 ms with Best-of-3 selection. Third, a hybrid pipeline combining complementary generators with SPICE-based ranking achieves 99.9% simulation validity using only 8 SPICE evaluations, 40 fewer than genetic algorithms.
Per-design quality remains below search baselines (5.48 vs. 7.48 reward). But ARCS is 1000 faster, and ARCS-seeded GA recovers 96.6% of cold-start quality with 49% fewer simulations. The two paradigms are complementary. Amortized circuit generation is not a replacement for search. It is a practical starting point.
Appendix A Per-Topology Results
Table XI presents hybrid-ranked results for all 32 evaluated topologies (50 samples each, bootstrap 95% CI), sorted by mean reward.
| Topology | Cat | Reward | 95% CI |
| current_mirror | B | 8.00 | [8.00, 8.00] |
| voltage_doubler | P | 7.96 | [7.96, 7.97] |
| shunt_regulator | R | 7.77 | [7.64, 7.85] |
| series_regulator | R | 7.59 | [7.41, 7.73] |
| sallen_key_lowpass | F | 7.49 | [7.36, 7.60] |
| common_emitter | B | 7.05 | [6.90, 7.20] |
| wien_bridge | O | 7.00 | [7.00, 7.00] |
| colpitts | O | 7.00 | [7.00, 7.00] |
| common_collector | B | 7.00 | [7.00, 7.00] |
| hartley | O | 7.00 | [7.00, 7.00] |
| inverting_amp | A | 6.81 | [6.76, 6.86] |
| phase_shift | O | 6.80 | [6.50, 7.00] |
| common_base | B | 6.67 | [6.56, 6.80] |
| differential_amp | A | 6.63 | [6.54, 6.73] |
| inv_summing_amp | A | 6.60 | [6.50, 6.68] |
| half_bridge | P | 6.52 | [6.31, 6.72] |
| cuk | P | 6.49 | [6.26, 6.72] |
| cascode | B | 6.42 | [6.31, 6.53] |
| charge_pump | P | 6.41 | [6.40, 6.42] |
| buck | P | 6.37 | [6.25, 6.49] |
| noninverting_amp | A | 6.30 | [6.14, 6.44] |
| push_pull | P | 6.04 | [5.77, 6.32] |
| boost | P | 6.01 | [5.85, 6.17] |
| transimpedance_amp | A | 6.00 | [6.00, 6.00] |
| twin_t_notch | F | 5.76 | [5.54, 5.98] |
| zeta_converter | P | 5.71 | [5.51, 5.93] |
| instrum_amp | A | 5.40 | [5.24, 5.58] |
| buck_boost | P | 5.34 | [5.22, 5.48] |
| sepic | P | 5.24 | [4.99, 5.50] |
| sk_highpass | F | 5.00 | [5.00, 5.00] |
| sk_bandpass | F | 4.99 | [4.91, 5.06] |
| state_var_filter | F | 4.45 | [4.30, 4.63] |
Current mirrors and voltage doublers achieve near-perfect scores (8.0) due to simple operating conditions. The weakest topologies, state variable filter (4.45) and Sallen-Key bandpass/highpass (5.0), reflect the difficulty of precise frequency control in multi-pole filters.